OpenStudy (anonymous):

A car is already at 40 km/h when the driver decides to increase his speed, he steps on the gas pedal for 6 seconds shows a speed of 100 km/h. What was the Acceleration ?

OpenStudy (anonymous):

acceleration = change of speed/change of time

OpenStudy (kc_kennylau):

Here is my list of unit lol distance=km (km h^0) speed=km/h (km h^-1) acceleration=km/h/h (km h^-2)